# Juding the Judges: A Systematic Investigation of Position Bias in Pairwise Comparative Assessments by LLMs

### v4

@@ Abstract
LLM-as-a-Judge presents a promising alternative to human evaluators across various tasks, but inherent biases, especially position bias — a tendency to favor solutions based on their position in the prompt — have compromised its effectiveness. Our study introduces a systematic framework to examine position bias in pairwise comparisons, focusing on repetition stability, position consistency, and preference fairness. This research significantly contributes to the field by introducing new concepts for understanding position bias and providing a multi-dimensional framework for evaluations. We conducted experiments with 12 LLM judges across MTBench and DevBench, covering 22 tasks and approximately 40 solution-generating models — candidates, resulting in over 100,000 evaluation instances. Our findings confirm that position bias in capable LLM judges is not due to random chances, along with notable variations observed across judges and tasks. Moreover, position bias is weakly influenced by the length of prompt components but significantly impacted by the quality gap between solutions. These insights can help optimize judge model selections, improve benchmark design, and inform future research on debiasing strategies, ultimately enhancing the reliability of LLM judges.
1 Introduction
@@
In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global interest and stimulating substantial research into their applications. Evaluating large language models (LLMs) has received increasing attention due to their advancing capabilities across diverse fields. While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations (Zeng et al., 2023; Karpinska et al., 2021). To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks. Despite a high level of agreement with human judgments(Zheng et al., 2024b; Li et al., 2024; Zhu et al., 2023b), inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
@@
Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than the content itself, as shown in Fig. 1. This bias is prevalent across all categories of LLM-as-a-Judge, including pointwise, listwise, and pairwise evaluations (Qin et al., 2024), for any score-based and relation-based settings (Li et al., 2023d). Consequently, the reliability of LLM judges becomes questionable when they exhibit position bias. However, addressing this issue is highly complicated due to the varying performance of LLM judges across different tasks (Khan et al., 2024; Chua et al., 2024). Moreover, the limited understanding of the bias itself has exacerbated the problem.
@@
Figure 1: Demonstration of position bias and repetition stability of LLM-as-a-Judge. When the positions of candidate model solutions are swapped in the prompt, position bias is exhibited if the LLM judge tends to favor the same position (e.g., always choosing {B}) rather than the same candidate model solution. We observe repetition instability if the same LLM judge gives different judgments across repetitions. Only when an LLM judge has high repetition stability can we validate that its position bias is not due to random variations.
Another key aspect of position bias, preference fairness, has also been under-explored by previous works. Most studies have primarily focused on position consistency to measure how frequently position bias occurs across a set of evaluation instances. However, the direction of biased preference (i.e., primacy or recency) is also insightful. Fig. 2 illustrates cases where a pair of judgments with swapped prompts become position consistent, inconsistent, primacy-preferred, or recency-preferred. We measure preference fairness by analyzing the distribution of these preference directions across a set of evaluations. Specifically, judgments are considered to have a fair preference if they are position consistent or if the preference directions are evenly distributed. In contrast, an LLM judge is considered to have an unfair preference if it consistently favors candidate model solution that appear either first or last in the prompt. Both fair and unfair judge models can provide valuable insights into model properties, applications, and strategies for improvement.
Figure 2: Demonstration of position consistent/inconsistent, primacy-preferred, and recency-preferred judgment pairs in pairwise comparative assessment, where the task is to choose the better candidate solution. A judgment is consistent if it selects the same winner across different prompt positions. Primacy/recency indicates a preference for the first/last candidate solution, respectively. Position inconsistent judgments suggest position bias and can be classified as primacy-preferred or recency-preferred.
To enhance the understanding of position bias and LLM-as-a-Judge, we propose a comprehensive framework designed to dissect and analyze position bias within the context of pairwise comparative assessment. We focus on this context because LLM judges demonstrate superior performance in making pairwise comparisons over other settings like pairwise scoring (Zheng et al., 2024b; Liusie et al., 2024), making it a practically preferable choice for investigations.
@@
2 Methods & Defintions
LLM-as-a-Judge can be classified into two categories: score-based and relation-based assessments. Our study focuses on relation-based assessment, specifically under the pairwise comparison setting, for several reasons. First, the pairwise paradigm is a popular and classic scenario that has influenced a wide range of fields (Qin et al., 2024). Second, as noted by Liu et al. (2024), LLM and human evaluations are more aligned in the context of pairwise comparison than any of the score-based assessments. Additionally, numerous studies have shown that pairwise comparative assessment outperforms other judging settings in terms of position consistency(Zheng et al., 2024b; Liusie et al., 2024). Finally, pairwise comparisons can be extended to more complicated frameworks of relation-based assessment, such as listwise comparisons, through some advanced ranking algorithms (Qin et al., 2024; Liu et al., 2024). Therefore, analyzing position bias in this context can provide a deeper understanding of its fundamental nature.
In the context of pairwise comparative assessments, LLM judges are prompted to choose the better candidate solution to the task question. To accommodate the possibility of a tie, various option modes are introduced. Two-Option mode restricts judges to choosing the better solution between two options, labeled {A} for the first and {B} for the second as shown in Fig. 1. Three-Option mode adds an additional choice, {C}, allowing judges to indicate a tie if neither solution is preferable. Four-Option mode further expands the choices, with {C} representing a both good tie and {D} a both bad tie. In our experiments, we employed Three-Option mode for MTBench (Zheng et al., 2024b) and Two-Option mode for DevBench (Li et al., 2024), aligning with their respective original frameworks. These option modes are explicitly specified in the system prompts to clearly guide the decision-making process of the LLM judges.
To set up our experiment, as shown in Fig. 1, the system prompt, option mode, task question, and solutions from two candidate models (original prompt) are presented to the LLM judges to select the winning solution. The experiment is conducted in a double-blind setting. The identities of the candidate models are hidden from the LLM judges, and the candidate models are unaware that their solutions will be compared to another model when answering the question. Then, the prompt with solutions in a swapped position (swapped prompt) is given to the same judge again, which results in a judgment pair. As shown in Fig. 2, if the LLM judge consistently favors the same solution regardless of the swapped position, it is considered position consistent with a fair preference. Conversely, if the LLM judge selects different winners, position bias is observed, with the preference direction being either primacy or recency.
@@ 2 Methods & Defintions
2.1 Evaluation Metrics
We validate the capable LLMs with a high repetition stability and then evaluate their position bias in terms of position consistency and preference fairness. The metrics are introduced as follows.
Repetition Stability ($RS$) evaluates the reliability of LLM judges when presented with identical queries multiple times. It is essential to determine whether the judgments of LLMs, and consequently the observations of position bias, stem from a consistent evaluation pattern or merely random variations. We measure this by calculating the percentage of the most frequent selections across multiple trials for each query, aggregated from all queries within each dataset. This metric is formalized as
$RC=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n_{j}}\max_{\forall k\in S}\left\{|C_{k}^{j}|\right\},\quad S=\{A,B,C,...\}$ (1)
@@ 2.1 Evaluation Metrics
Position Consistency ($PC$) quantifies how frequently a judge model prefers the same solution after the order of solutions is swapped. It is calculated as the ratio of consistent evaluation pairs to the total number of valid evaluations, where a pair is deemed consistent if the judge model prefers the same winning solution. Formally, it is calculated as
$PC=\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})\in V\}},V=\{(A,B),(B,A),(C,C),...\}$ (2)
@@
Preference Fairness ($PF$) is another crucial characteristic of position bias measuring the extent to which judge models favor certain solution positions. In the pairwise comparative setup, an LLM judge may exhibit a preference for either primacy or recency. These terms replace the more verbose “preference for the first/second position” used in previous studies, ensuring clarity and generalization for future research. The examples of such preferences are demonstrated in Fig. 2.
Position consistent judgments are fair in preference. The measurements become more complex for position inconsistent cases. Previous studies proposed two common ways to measure the preference fairness. A straightforward way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$). The counts are then normalized by the total number of prompt pairs (Zheng et al., 2024b; Zhu et al., 2023b). However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
Alternatively, instead of normalizing over the complete dataset, studies like (Li et al., 2023c; Liusie et al., 2024) treat position inconsistent evaluation instances independently. They calculate the percentages of primacy-preferred and recency-preferred judgment pairs relative to the total number of position inconsistent pairs. We denote these as inconsistent primacy rates ($ipr$) and inconsistent recency rates ($irr$), where $ipr+irr=1$. However, this approach is problematic because it overlooks the fact that “position consistent judgments are also position fair,” which can lead to overly penalizing highly consistent LLM-judges. For instance, if only one out of a hundred pairs is inconsistent, this method would classify it as 100% primacy-preferred or recency-preferred. However, this outcome should be distinguished from cases where all one hundred pairs are inconsistent and biased toward one position, which would also result in 100% preference unfairness.
@@
$PF=\frac{PF_{raw}-S^{-}_{min}}{S^{+}_{max}-S^{-}_{min}}\times 2-1,PF_{raw}=(rcn\times irr)-(pcn\times ipr)$ (3)
where $S^{-}_{min}$ and $S^{+}_{max}$ are the minimum and maximum achievable $PF_{raw}$ scores for each judge on each task, respectively. This min-max scaling technique ensures comparability across datasets by accounting for the range of achievable scores and centering the scale around zero. Compared to prior methods, the use of $rcn$, $pcn$ weighted by $irr$ and $ipr$, and scaling across judges and tasks implicitly accounts for the position consistent judgments. Another advantage of this approach is that it quantifies preference fairness with a single score, rather than the two separate scores used in previous studies. The $PF$ score is interpreted as follows:
$PF=\begin{cases}1,&\text{if }PC=0\text{ and entirely recency-preferred}\\ x\in(0,1),&\text{Recency-preferred}\\ 0,&\text{Preference Fair}\\ x\in(-1,0),&\text{Primacy-preferred}\\ -1,&\text{if }PC=0\text{ and entirely primacy-preferred}\end{cases}$
@@
2.2 Factors affecting position bias
Table 1: Categorization of the factors that influence position bias, where ‘Judge’ refers to the LLM-judge and ‘Candidate’ refers to the candidate model whose solutions to the task-specific questions are evaluated by the judges. Factors that have a significant impact on position bias are highlighted with an asterisk (*) by bidirectional stepwise regression tests and marked red by empirical findings. ‘Task Input’ refers to the question itself, while ‘Task Output’ refers to the candidate model’s solutions to the question, making it both a Candidate-level and Task-level factor. ‘Prompt’ encompasses the complete query presented to the judges, including Task Input, Task Output, and system prompts.
@@ 2.2 Factors affecting position bias
Task Output Length $\times$ $\checkmark$* $\checkmark$*
@@
Among the influencing factors we identified, we chose “familial property” for Judge-level factors instead of model sizes and training specifics, which are often proprietary and not publicly accessible for the models involved in our experiments. The model family grouping reflects similar model sizes and training specifics of the model members. The familial categories of the models used in our studies are (1) GPT, (2) Claude, and (3) Gemini allowing for straightforward grouping by company and version. However, we also observed that all GPT-4 and later models, along with Claude-3.5-Sonnet, Claude-3-Opus, and Gemini-1.5-pro, form a group of the “most capable” models sharing similar judging capabilities. More details and discussions about the familial property can be found in Appendix Sec. C, where we examined the LLM mutual agreements to study the similarities and differences across families.
Answer quality gap, based on our findings, is highly impactful yet under-explored in prior research. We define the quality of a candidate’s answer/solution by how effectively it addresses the question. Consequently, the answer quality gap refers to the disparity in quality between the solutions from two candidate models to the same question and is considered the Candidate-level (i.e., solution provider) factor. Ideally, when a reliable LLM judge is presented with a question and corresponding solution pairs, it would prefer the higher-quality solution, where the corresponding candidate is denoted as the winner selected by the judge. Following this assumption, we measure the answer quality gap by the win rates between two candidates on a set of tasks and questions. We regard a tie judgment as ‘half-win’ for both candidates when calculating its win rate over another, depending on option mode. If position bias occurs, the winner may be inconsistent when the order of candidate solutions is swapped in the query. Therefore, we categorize the LLM judgments into three groups: cases where the same winner is consistently chosen in both the original and swapped queries (termed “consistent wins”), cases where both responses are deemed equal (termed “consistent ties”), and cases where different winners are selected after the solutions are swapped. We denote these counts as the number of consistent wins ($C_{w}$), consistent ties ($C_{t}$), and inconsistent judgment pairs ($C_{I}$), respectively. Following Zheng et al. (2024b), we count inconsistent judgment pairs also as ties, or ‘half-wins’, to ensure calculations are feasible in all cases for the subsequent computations.
@@
Inspired by Zheng et al. (2024b), we define the overall win rate ($owr$) of a model’s solution over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of judgment pairs. Then the answer quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-0.5|$, where 0.5 is the overall win rate when we only have ties. In contrast to using only consistent win rate (calculated as $\frac{C_{w}}{n_{c}}$, where $n_{c}$ is the number of position consistent judgment pairs) to quantify $\delta_{q}$ (Zheng et al., 2024b; Li et al., 2023b; Raina et al., 2024), the adoption of overall win rate incorporates all data points and captures the “comparable quality” cases, where responses in similar quality might lead to position biased judgments, a scenario that the consistent win rate might overlook. More discussions can be found in Appendix. Sec. D.
3 Experiments
3.1 Experiment Settings

### v5

# Juding the Judges: A Systematic Investigation of Position Bias in Pairwise Comparative LLM-as-a-Judge

@@ Abstract
LLM-as-a-Judge offers a promising alternative to human judges across various tasks, yet inherent biases, particularly position bias—a consistent preference for answers based on their position in the prompt—compromise its effectiveness. Our study proposes a systematic framework to investigate the position bias of LLM judges in the pairwise comparative scenario in terms of repetition stability, position consistency, and preference fairness. In pairwise comparative settings, LLM judges are prompted to choose the better candidate’s answer to the task question, where the task question and a pair of candidate answers are provided. Then the prompt with swapped order of candidate answers is given to the same judge again. If the judge prefers the same winner, it is denoted position consistent; otherwise, if different winners are selected after swapped settings, position bias is observed.
To comprehensively understand the position bias of LLM judges, our experiments cover 12 judges across MTBench and DevBench with 22 tasks and around 40 answer-generating models, resulting in over 100,000 evaluation instances. Further exploration of the Judge-level, Candidate-level, and Task-level impactful factors on position bias enhances the understanding of this bias. Through comprehensive investigations, we validate that the position bias of capable LLM judges is not merely random variations. We also observe significant variance of position bias across judges and tasks, highlighting trade-offs of positional consistency, preference fairness, and practicality when selecting task-specific judges. Among impactful factors, position bias is largely influenced by the quality gap of given candidate answer pairs for evaluation. However, the lengths of the prompt components, including input questions, candidate answers, and the complete prompt, have minimal impact on position bias. In particular, explorations of answer quality gap and LLM agreements not only help explain length/verbosity bias and self-enhancement bias from previous studies, but also provide evidence for distinguishing difficult-to-evaluate instances that benefit future benchmark designs. These insights guide the selection of optimal judge models, enhance benchmark design, and lay the foundation for future research into effective debiasing strategies, ultimately improving the reliability of LLM evaluators for more robust, unbiased evaluation systems.
Provocations
Despite advancements in LLM-as-a-Judge and prevalent attention on position bias, the understanding of such bias is still limited and requires a more systematic evaluation. Enhancing the understanding of the impactful factors on position bias could inspire more effective mitigation strategies, laying the foundation for more reliable, robust, and unbiased evaluation systems. To this end, our study investigates aspects that are underexplored by prior works to offer a systematic evaluation framework for the position bias of LLM judges.
Repetition Stability
Validating LLM judges are repetition stable is a primary challenge that is often ignored by previous studies. They trusted the one-time generated judgments of the LLMs and conducted further analysis, implicitly assuming that the judgments represent the judge’s capabilities. However, the judgment choices and consequently observed position bias may be due to random variations. In other words, without verifying that LLM judges are stable in their choices across multiple trials, it is uncertain whether position bias stems from positional shift or is merely randomness. To address this gap, our study utilizes repetition stability to validate capable LLMs before further investigations on position bias.
Preference Fairness
While position bias is prevalent across most tasks and widely studied by literature, the main focus is on position consistency, as it is the most straightforward measurement of position bias. However, the preference directions when position bias occurs are largely underexplored by prior works. In this study, we show that previous measurements are either sensitive to the dataset size or overly penalizing highly consistent judges. To address this problem, we propose a more granular and scalable measurement that combines the strengths of both methods to assess preference fairness. We further stress the trade-off between position consistency, preference fairness, and practicality when choosing the task-specif judges, as position bias largely varies by task and judge.
Answer Quality Gap
From intuition, the difficulty of judging a pair of candidate answers (i.e., choosing the winner that better answers the task question) largely depends on the disparities of the answer pairs in quality. However, we are the first study to quantify the answer quality gap and investigate its correlation with position bias. Our investigations verify that as the answer quality gap enlarges, implying one candidate’s answer is significantly better than the other, the judges become more constant and fair. The findings not only benefit the interpretation of other LLM judge biases like verbosity/length bias and self-enhancement bias, but also benefit future benchmark designs for more social goods.
To summarize, this study proposes new metrics and studies more impacting factors on position bias that are underexplored by prior studies, offering a more comprehensive and systematic framework for understanding and evaluating the position bias of LLM judges. We stress the need to have a deeper understanding of the complex bias before moving to solutions. More detailed analysis, insights, findings, and contributions can be found in the Appendix.
[…]
[appendices]
Appendix Table of Contents
[…]
[appendices]0
Appendix
Appendix A Introduction
@@
In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global attention and stimulating substantial research into their applications. Evaluating large language models (LLMs) has received increasing attention due to their advancing capabilities across diverse fields. While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations [1, 2]. To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks. Despite a high level of agreement with human judgments[3, 4, 5], inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
@@
Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than the content itself, as shown in Fig. 1. This bias is prevalent across all categories of LLM-as-a-Judge, including pointwise, listwise, and pairwise evaluations [6], for any score-based and relation-based settings [7]. Consequently, the reliability of LLM judges becomes questionable when they exhibit position bias. However, addressing this issue is highly complicated due to the varying performance of LLM judges across different tasks [8, 9]. Moreover, the limited understanding of the bias itself has exacerbated the problem.
@@
Figure 1: Demonstration of position bias and repetition stability of LLM-as-a-Judge. When the positions of model responses are swapped in the prompt, position bias is exhibited if the LLM judge tends to favor the same position (e.g., always choosing {B}) rather than the same response. We observe repetition instability if the same LLM judge gives different judgments across repetitions. Only when an LLM judge has high repetition stability can we validate that its position bias is not due to random variations.
Another key aspect of position bias, preference fairness, has also been under-explored by previous works. Most studies have primarily focused on positional consistency to measure how frequently position bias occurs across a set of evaluation instances. However, the direction of biased preference (i.e., primacy or recency) is also insightful. Fig. 2 illustrates cases where a pair of judgments with swapped prompts become positionally consistent, inconsistent, primacy-preferred, or recency-preferred. We measure preference fairness by analyzing the distribution of these preference directions across a set of evaluations. Specifically, judgments are considered to have a fair preference if they are positionally consistent or if the preference directions are evenly distributed. In contrast, an LLM judge is considered to have an unfair preference if it consistently favors responses that appear either first or last in the prompt. Both fair and unfair judge models can provide valuable insights into model properties, applications, and strategies for improvement.
Figure 2: Demonstration of positionally consistent/inconsistent, primacy-preferred, and recency-preferred judgment pairs in pairwise comparative assessment, where the task is to choose the better response. A judgment is consistent if it selects the same winner across different prompt positions. Primacy/recency indicates a preference for the first/last model response, respectively. Positionally inconsistent judgments suggest position bias and can be classified as primacy-preferred or recency-preferred.
To enhance the understanding of position bias and LLM-as-a-Judge, we propose a comprehensive framework designed to dissect and analyze position bias within the context of pairwise comparative assessment. We focus on this context because LLM judges demonstrate superior performance in making pairwise comparisons over other settings like pairwise scoring [3, 10], making it a practically preferable choice for investigations.
@@
Appendix B Methods & Defintions
LLM-as-a-Judge can be classified into two categories: score-based and relation-based assessments. Our study focuses on relation-based assessment, specifically under the pairwise comparison setting, for several reasons. First, the pairwise paradigm is a popular and classic scenario that has influenced a wide range of fields [6]. Second, as noted by [11], LLM and human evaluations are more aligned under the context of pairwise comparison than any of the score-based assessments. Additionally, numerous studies have shown that pairwise comparative assessment outperforms other judging settings in terms of positional consistency[3, 10]. Finally, pairwise comparisons can be extended to more complicated frameworks of relation-based assessment, such as listwise comparisons, by some advanced ranking algorithms [6, 11]. Therefore, analyzing position bias in this context can provide a deeper understanding of its fundamental nature.
In the context of pairwise comparative assessments, LLM-judges are prompted to choose the better response that answers the question. To accommodate the possibility of a tie, various option modes are introduced. The Two-Option mode restricts judges to choosing the better response between two options, labeled {A} for the first and {B} for the second as shown in Fig. 1. The Three-Option mode adds an additional choice, {C}, allowing judges to indicate a tie if neither option is preferable. The Four-Option mode further expands the choices, with {C} representing a both good tie and {D} a both bad tie. In our experiments, we employ the Three-Option mode for MTBench [3] and the Two-Option mode for DevBench [4], aligning with their respective original frameworks. These option modes are explicitly specified in the system prompts to clearly guide the decision-making process of the LLM-judges.
To setup our experiment, as shown in Fig. 1, the system prompt, option mode, task question, answers/responses from the two candidate models (original prompt) are presented to the LLM-judges to select the winning answer. The experiment is conducted in a double-blind setting. The identities of the candidate models are hidden from the LLM-judges, and the candidate models are unaware that their responses will be compared to another model when answering the question. Then, the prompt with answers in a swapped position (swapped prompt) is given to the same judge again, which results in a choice pair. As shown in Fig. 2, if the choice pair consistently favors the same answer regardless of the swapped position, it is considered positionally consistent with a fair preference. Conversely, if the LLM-judge selects different winners, position bias is observed, with the preference direction being either primacy or recency.
@@ Appendix B Methods & Defintions
B.1 Evaluation Metrics
Repetition Stability ($RS$) evaluates the reliability of LLM judges when presented with identical queries multiple times. It is essential to determine whether the choices of LLM judges, and consequently the observations of position bias, stem from a consistent evaluation pattern or merely random variations. We measure this by calculating the percentage of the most frequent selections across multiple trials for each query, aggregated from all queries within each dataset. This metric is formalized as
$RC=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n_{j}}\max_{\forall k\in S}\left\{|C_{k}^{j}|\right\},\quad S=\{A,B,C,D\}$
@@ B.1 Evaluation Metrics
Position Consistency ($PC$) quantifies how frequently a judge model selects the same response after the order of responses is swapped. It is calculated as the ratio of consistent evaluation pairs to the total number of valid evaluations, where a pair is deemed consistent if the model selects the same winner. Formally, it is calculated as
$PC=\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})\in V\}}$
@@
Preference Fairness ($PF$) is another crucial characteristic of position bias measuring the extent to which judge models favor certain response positions. In the pairwise comparative setup, an LLM judge may exhibit a preference for either primacy or recency. These terms replace the more verbose “preference for the first/second position" used in previous studies, ensuring clarity and generalization for future research. The examples of such preferences are demonstrated in Fig. 2.
Positionally consistent judgments are fair in preference. The measurements become more complex for positionally inconsistent cases. Previous studies proposed two common ways to measure the preference fairness. A straightforward way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$). The counts are then normalized by the total number of prompt pairs [3, 5]. However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
Alternatively, instead of normalizing over the complete dataset, studies like [12, 10] treat positionally inconsistent evaluation instances independently. They calculate the percentages of primacy-preferred and recency-preferred judgment pairs relative to the total number of positionally inconsistent pairs. We denote these as inconsistent primacy rates ($ipr$) and inconsistent recency rates ($irr$), where $ipr+irr=1$. However, this approach is problematic because it overlooks the fact that “positionally consistent judgments are also positionally fair," which can lead to overly penalizing highly consistent LLM-judges. For instance, if only one out of a hundred pairs is inconsistent, this method would classify it as 100% primacy-preferred or recency-preferred. However, this outcome should be distinguished from cases where all one hundred pairs are inconsistent and biased toward one position, which would also result in 100% preference unfairness.
@@
$PF=\frac{PF_{raw}-S^{-}_{min}}{S^{+}_{max}-S^{-}_{min}}\times 2-1,PF_{raw}=(rcn\times irr)-(pcn\times ipr)$
where $S^{-}_{min}$ and $S^{+}_{max}$ are the minimum and maximum achievable $PF_{raw}$ scores for each judge on each task, respectively. This min-max scaling technique ensures comparability across datasets by accounting for the range of achievable scores and centering the scale around zero. Compared to prior methods, the use of $rcn$, $pcn$ weighted by $irr$ and $ipr$, and scaling across judges and tasks implicitly accounts for the positionally consistent judgments. Another advantage of this approach is that it quantifies positional preference with a single score, rather than the two separate scores used in previous studies. The $PF$ score is interpreted as follows:
$PF=\begin{cases}1,&\text{if }PC=0\text{ and entirely recency-preferred}\\ x\in(0,1),&\text{Recency-preferred}\\ 0,&\text{Positionally Fair}\\ x\in(-1,0),&\text{Primacy-preferred}\\ -1,&\text{if }PC=0\text{ and entirely primacy-preferred}\end{cases}$
@@
B.2 Factors affecting position bias
Table 1: Categorization of the factors that influence position bias, where ‘Judge’ refers to the LLM-judge and ‘Candidate’ refers to the candidate model whose responses to the task-specific questions are evaluated by the judges. An asterisk (*) highlights the factors with significant impact on position bias that we found through experiments. ‘Task Input’ refers to the question itself, while ‘Task Output’ refers to the candidate model’s answers to the question, making it both a Candidate-level and Task-level factor. ‘Prompt’ encompasses the complete query presented to the judges, including Task Input, Task Output, and system prompts.
@@ B.2 Factors affecting position bias
Task Output Length $\times$ $\checkmark$ $\checkmark$
@@
Among the influencing factors we identified, we chose “familial property” instead of model sizes and training specifics, which are often proprietary and not publicly accessible for the models involved in our experiments. The model family grouping reflects similar model sizes and training specifics of the model members. The familial categories of the models used in our studies are (1) GPT, (2) Claude, and (3) Gemini allowing for straightforward grouping by company and version. However, we also observed that all GPT-4 and later models, along with Claude-3.5-Sonnet, Claude-3-Opus, and Gemini-1.5-pro, form a group of the “most capable" models sharing similar judging capabilities. More details and discussions about the familial property can be found in Appendix Sec. H, where we examined the LLM mutual agreements to study the similarities and differences across families.
Answer quality gap, based on our findings, is a highly impactful yet under-explored factor in prior research. We define the quality of a model’s answer by how effectively it addresses the question. Consequently, the answer quality gap refers to the disparity in quality between two candidate model answers to the same question and is considered a Candidate-level (i.e., response provider) factor. Ideally, when a reliable LLM judge is presented with a question and corresponding answer pairs, it would prefer the higher-quality answer, where the corresponding candidate is denoted as the winner selected by the judge.
Depending on option mode, we regard a tie judgment as‘half-win’ for both candidates when calculating its win rate over another.
Then, the win rates between two candidates on a set of task questions can reflect the answer quality gap between them on the task.
However, if position bias is exhibited, the winner may be inconsistent when the order of candidate responses is swapped in the query. Therefore, we observe the LLM judgments where the absolute winner is consistently chosen in both original and swapped queries (termed “consistent wins”), both responses are deemed equal (termed “consistent ties”), and different winners are selected after candidate responses are swapped in the query. We denote these counts as the number of consistent wins ($C_{w}$), consistent ties ($C_{t}$), and inconsistent choice pairs ($C_{I}$), respectively. Following [3], we count inconsistent choice pairs as ties, or ‘half-wins’, to ensure calculations are feasible in all cases for the subsequent computations.
@@
Inspired by [3], we define the overall win rate ($owr$) of a model’s response over the other as: $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of choice pairs. Then the answer quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-0.5|$, where 0.5 is the overall win rate when we only have ties. In contrary to using consistent win rate by [3] (calculated as $\frac{C_{w}}{n_{C}}$, where $n_{C}$ is the number of positionally consistent judgment pairs) to quantify $\delta_{q}$, the adoption of overall win rate incorporates all data points and captures the “comparable quality” scenario, where responses in similar quality might lead to positionally biased judgments, a scenario that the consistent win rate might overlook. More discussions can be found in Appendix. Sec. I.
Appendix C Experiments
C.1 Experiment Settings